OpenStudy (anonymous):

HELP PLEASE. What is the value of x? 3√9+x=15

OpenStudy (anonymous):

let's start with \[\sqrt9\] what does that equal?

OpenStudy (anonymous):

make a guess

OpenStudy (anonymous):

3

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

\[3\cdot3+x=15\]

OpenStudy (anonymous):

what is 3 times 3?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok so we now have \[9+x=15\] how do we get the 9 to the other side of the equal sign?

OpenStudy (anonymous):

make a guess

OpenStudy (anonymous):

i really don't know :/ i have been out of math for 3 years now. :(

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

we would subtract 9 from both sides to get x by itself so \[\cancel{9-9}+x=15-9\]

OpenStudy (anonymous):

Yes! Good job!

OpenStudy (anonymous):

THANK YOUUUU!!!!

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

medal please? =D

OpenStudy (anonymous):

did i do the medal right?? I've never used this site before lol

OpenStudy (anonymous):

I got it. Thank you!

OpenStudy (anonymous):

you're welcome. thank you so much as well!!

OpenStudy (anonymous):

Glad to help!